Question:
A mixture of 20% acid and 50% acid must be mixed together to obtain 100mL of 44% acid solution. How much of each acid will be required?
Answer:
Let x represent amount (in mL) of 20% acid required
Let y represent amount (in mL) of 50% acid required
100 = x + y
44 = 0.2x + 0.5y
I got the right answer
x = 20
y = 80
I just don't understand why 44 isn't changed into 0.44 and the other numbers are. Would you mind explaining?

Answer by MathTherapy(10552)


Question:
A mixture of 20% acid and 50% acid must be mixed together to obtain 100mL of 44% acid solution. How much of each acid will be required?
Answer:
Let x represent amount (in mL) of 20% acid required
Let y represent amount (in mL) of 50% acid required
100 = x + y
44 = 0.2x + 0.5y
I got the right answer
x = 20
y = 80
I just don't understand why 44 isn't changed into 0.44 and the other numbers are. Would you mind explaining?

x + y = 100 ----- eq (i)
0.2x + 0.5y = .44(100) ------ 0.2x + 0.5y = 44

Do you now see how 44 was derived?

Send comments and “thank-yous” to “D” at MathMadEzy@aol.com